There are 112 male and 78 females working at a hospital. What percentage of the hospitals workforce is female?

Found 2 solutions by JBarnum, nyc_function:
Answer by JBarnum(2146)

112+78=total workforce=190
so 78/190=.410526=41.0526%

Answer by nyc_function(2741)

Let p = percent of hospital force that is female
p = number of females divided by number of men
p = 78/112
p = 0.69 or 69%